# Seoul 121-742, South Korea

### v1

@@ 4.3 Theoretical Challenges and Future Directions
The most significant limitation concerns the treatment of strong phase transitions with $\alpha>1$. Current numerical simulations of bubble dynamics and gravitational wave production are calibrated primarily for weak-to-moderate transitions, and their extrapolation to the vacuum domination regime introduces systematic uncertainties of order unity in the predicted amplitude. Improved simulations incorporating relativistic magnetohydrodynamics and non-linear gravitational effects are essential for robust predictions in this parameter range. The most significant limitation concerns the treatment of strong phase transitions with $\alpha>1$. Current numerical simulations of bubble dynamics and gravitational wave production are calibrated primarily for weak-to-moderate transitions, and their extrapolation to the vacuum domination regime introduces systematic uncertainties of order unity in the predicted amplitude. Improved simulations incorporating relativistic magnetohydrodynamics and non-linear gravitational effects are essential for robust predictions in this parameter range.
The calculation of bubble wall velocity represents another crucial source of uncertainty. While we employ state-of-the-art tools like WallGo Ekstedt:2024fyq for selected benchmarks (specifically for the $SU(2)\rightarrow\mathds{1}$ case), the systematic scan over parameter space relies on approximate analytical expressions that may not capture the full complexity of wall dynamics in flat potential scenarios. The interplay between friction from particle interactions and the driving pressure from the effective potential requires more sophisticated treatment, particularly near the transition between deflagration and detonation modes.
@@ 5 Conclusions
The detection of gravitational waves offers a promising new avenue to probe physics beyond the Standard Model, particularly in regimes inaccessible to current collider and laboratory experiments. In this work, we have studied the gravitational wave signatures of a first-order phase transition associated with the spontaneous breaking of lepton number symmetry, which dynamically generates Majorana masses for right-handed neutrinos. This mechanism is strongly motivated by the compelling evidence for neutrino masses and mixing, and it naturally arises in minimal seesaw frameworks capable of explaining the observed matter–antimatter asymmetry via leptogenesis. Our flat potential approach provides a particularly natural and robust framework for realizing strong first-order transitions without requiring fine-tuned scalar couplings or exotic particle content.
@@ Appendix B Further contributions to the thermal potential
This procedure remains valid in the presence of higher-dimensional operators like $\phi^{6}$, provided the temperature remains below the EFT cutoff ($T\ll m_{X}$) and the dimension-six coupling $\lambda/m_{X}^{2}$ is perturbative. In such cases, the dominant thermal correction arising from the $\phi^{6}$ operator scales as $T^{4}/m_{X}^{2}$, and can be included consistently in the thermal mass entering the Daisy resummation. For the $SU(2)$ cases the derivation is similar and yields exactly the same result.
@@ Appendix C Formulas for GW density
As is well established, the density of gravitational waves (GW) is governed by the following parameters: the nucleation temperature $T_{n}$, the strength parameter $\alpha$, the inverse duration of the phase transition $\beta$, the sound speed $c_{s}$, the bubble wall velocity $v_{w}$, as well as the efficiency factors $k_{\nu}$ and $\epsilon$. In terms of these quantities, one may employ the standard formulas for the contributions from sound waves and turbulence, as derived in Refs. Hindmarsh:2013xza ; Hindmarsh:2015qta and Caprini:2009yp ; Binetruy:2012ze , respectively. The contribution from sound waves is given by
$\Omega_{\rm sw}h^{2}(f)=2.65\times 10^{-6}(H_{*}\tau_{\rm sw})\left(\frac{H}{\beta}\right)\,v_{b}\left(\frac{\kappa_{\nu}\alpha}{1+\alpha}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-\frac{1}{3}}\left(\frac{f}{f_{\rm sw}}\right)^{3}\left(\frac{7}{4+3\left(f/f_{\rm sw}\right)^{2}}\right)^{7/2},$
where the factor $\tau_{sw}=\rm{min}\left[\frac{1}{H_{*}},\frac{R_{*}}{\bar{U}_{f}}\right]$ is the time scale of the duration of the phase Ellis:2019oqb . This quantity can be either $1/H_{*}$ or ${R_{*}}/{\bar{U}_{f}}$, where $H_{*}R_{*}=v_{b}(8\pi)^{1/3}(\beta/H)^{-1}$. The root-mean-square (RMS) fluid velocity can be approximated as
$\bar{U}_{f}^{2}~\approx~\frac{3}{4}\,\left(\frac{\kappa_{\nu}\alpha}{1+\alpha}\right)$
The MHD turbulence in the plasma is the sub-leading source of GW signals, whose energy density is given by
$\Omega_{t}h^{2}\,(f)=3.35\times 10^{-4}\left(\frac{\beta}{H}\right)^{-1}\left(\frac{\epsilon\,\kappa_{\nu}\alpha}{1+\alpha}\right)^{\frac{3}{2}}\left(\frac{g_{*}}{100}\right)^{-\frac{1}{3}}\,v_{b}\,\frac{\left(f/f_{\rm turb}\right)^{3}\left(1+f/f_{\rm turb}\right)^{-\frac{11}{3}}}{\left[1+8\pi\frac{f}{h_{*}}\right]}$
@@
$h_{*}=16.5\frac{T_{*}}{10^{8}\,\rm{GeV}}\left(\frac{g_{*}}{100}\right)^{1/6}\,\rm{Hz}.$ (112)
The work Caprini:2024gyk proposed updated templates, in the weak, medium, and strong ($\alpha\sim 0.5$) regimes. This last one for the first time. The basic result is that for weak FOPT we can trust the previous literature, while for a weak and strong FOPT, there is a modification of the assumed linear growth with the source duration.

### v2

@@ 4 Discussion
4.1 Connection to Neutrino Masses
The scope of this work is not to present a complete model of neutrino masses but to show that it is possible within the context of a theory of neutrino mass generation to produce a potential of the type that we have been describing and such that the effective mass $M_{R}$ can be associated to the see-saw mechanism. To this end, in the context of Type I seesaw, we have the well-known relation
$m_{\nu}\simeq\frac{y^{2}v^{2}}{M_{R}}.$ (57)
We note that, unlike the conventional Type-I seesaw realization where the Standard Model Higgs vacuum expectation value explicitly enters the neutrino mass formula, in our setup the dominant contribution to the Majorana mass scale arises from the vacuum expectation value of the scalar field responsible for spontaneous lepton-number breaking. As a result, the seesaw suppression is controlled by the lepton-number breaking scale rather than directly by the electroweak Higgs vev.
If we want to reproduce the mass of the atmospheric neutrinos, we need $m_{\nu}\sim 0.05\penalty 10000\ {\rm eV}$ ParticleDataGroup:2024cfk . In Tab. 4, we present representative benchmark points, covering the scales $10^{7}$ GeV to $\times 10^{10}$ GeV that can satisfy the relation Eq. (57), for one family of fermions within the model described in Section 2.3.
$v$ [GeV] $y$ $M_{R}$ [GeV]
$3.5\times 10^{5}$ $6.39\times 10^{-8}$ $1\times 10^{7}$
$1.1\times 10^{6}$ $6.42\times 10^{-8}$ $1\times 10^{8}$
$3.6\times 10^{6}$ $6.2\times 10^{-8}$ $1\times 10^{9}$
$0\times 10^{6} 8.8$$8\times 10^{-8}$ $1\times 10^{10}$
Table 4: Representative benchmark points that satisfy $m_{\nu}=0.05\,$eV.
The Yukawa couplings displayed in Table 4 are correspondingly small. While such values may appear to deviate from the usual naturalness motivation of the seesaw mechanism, they are technically natural in the ’t Hooft sense, since setting them to zero enhances the symmetry of the theory. In this work, we adopt these values as a consistent and minimal choice that allows us to focus on the phase transition dynamics and the resulting gravitational-wave signals.
In this section, we would also like to connect the neutrino scale to strong first-order phase transitions, and therefore, we focus on two benchmark points that enter this regime in the observable LISA region. In Fig. 13 we present the contour plots corresponding to lines 2 and 3 of Tab. 4.
Figure 13: For the $SU(2)\rightarrow U(1)$ case, we present two examples. Left: Second row of Tab. 4 that is for a scale of $m_{X}=10^{8}$ GeV, and Right: Fourth row of Tab. 4 that corresponds to a scale of $m_{X}=10^{10}$ GeV.
As expected, since the value of the Yukawa coupling is so small, it does not differ greatly from the case of $y=0$ and $n_{f}=0$, Fig. 4(b). This can be observed by comparing Fig. 4(b) (b) to Fig. 13.
If thermal inflation dilutes pre-existing asymmetry, leptogenesis must occur after the phase transition. This constrains $T_{R}>10^{9}$ GeV for standard thermal leptogenesis, which may conflict with lower $m_{X}$ values, but in this case we have scenarios for which $T_{R}>10^{9}$ GeV.
@@ 4.3 Multi-Messenger Observational Strategy
4.4 Cosmological History: Thermal Inflation Dynamics in Benchmark Scenarios
The number of e-foldings in thermal inflation can be quantified by noting that the era starts when $T\sim V_{0}^{1/4}\sim(m_{0}m_{X})^{1/2}$ and ends when $T\sim m_{0}$ Lyth:1995ka ; Barreiro:1996dx , therefore
$N\sim\frac{1}{2}\ln{\left(\frac{m_{X}}{m_{0}}\right)}.$ (58)
Hence, we have $N=3.4,4.6,5.8,6.9$ respectively for $m_{X}=10^{7},10^{8},10^{9},10^{10}$ GeV and $m_{0}=10^{4}$ GeV. Note that this estimation is independent of the value of $\alpha$.
Regarding scalar and tensor perturbations generated during this brief inflationary period, both are negligible compared to the primordial fluctuations produced during the earlier high-scale inflationary epoch. The amplitude of scalar perturbations generated during an inflationary phase scales as $\Delta_{\mathcal{R}}^{2}\sim(H/\dot{\phi})^{2}$, while tensor perturbations scale as $\Delta_{h}^{2}\sim(H/M_{\rm Pl})^{2}$. For the phase transition considered here, with $H\sim H_{*}\sim 10^{-16}M_{\rm Pl}$ (corresponding to $T_{*}\sim 10^{6}$ GeV), these contributions are suppressed by many orders of magnitude relative to the primordial signals and are therefore observationally negligible, independently of the precise duration of this vacuum-dominated phase. The primary observational consequence of this vacuum-dominated phase is the enhanced redshift of the gravitational wave signal from the phase transition itself, as discussed in Section 3.3.
@@ 4.5 Theoretical Challenges and Future Directions
The most significant limitation concerns the treatment of strong phase transitions with $\alpha>1$. Current numerical simulations of bubble dynamics and gravitational wave production are calibrated primarily for weak-to-moderate transitions, and their extrapolation to the vacuum domination regime introduces systematic uncertainties of order unity in the predicted amplitude. Improved simulations incorporating relativistic magnetohydrodynamics and non-linear gravitational effects are essential for robust predictions in this parameter range.
The calculation of bubble wall velocity represents another crucial source of uncertainty. While we employ state-of-the-art tools like WallGo Ekstedt:2024fyq ; vandeVis:2025plm for selected benchmarks (specifically for the $SU(2)\rightarrow\mathds{1}$ case), the systematic scan over parameter space relies on approximate analytical expressions that may not capture the full complexity of wall dynamics in flat potential scenarios. The interplay between friction from particle interactions and the driving pressure from the effective potential requires more sophisticated treatment, particularly near the transition between deflagration and detonation modes.
@@ 5 Conclusions
The detection of gravitational waves offers a promising new avenue to probe physics beyond the Standard Model, particularly in regimes inaccessible to current collider and laboratory experiments. In this work, we have studied the gravitational wave signatures of a first-order phase transition associated with the spontaneous breaking of lepton number symmetry, which dynamically generates Majorana masses for right-handed neutrinos. This mechanism is strongly motivated by the compelling evidence for neutrino masses and mixing, and it naturally arises in minimal seesaw frameworks capable of explaining the observed matter–antimatter asymmetry via leptogenesis. Our flat potential approach provides a particularly natural and robust framework for realizing strong first-order transitions without requiring fine-tuned scalar couplings or exotic particle content. A striking outcome of the analysis is the emergence of a nontrivial scaling of the potential barrier, $V\propto m_{0}^{3}m_{X}$, revealing a distinctive and unexpected enhancement mechanism.
@@ Appendix B Further contributions to the thermal potential
This procedure remains valid in the presence of higher-dimensional operators like $\phi^{6}$, provided the temperature remains below the EFT cutoff ($T\ll m_{X}$) and the dimension-six coupling $\lambda/m_{X}^{2}$ is perturbative. In such cases, the dominant thermal correction arising from the $\phi^{6}$ operator scales as $T^{4}/m_{X}^{2}$, and can be included consistently in the thermal mass entering the Daisy resummation. For the $SU(2)$ cases, the derivation is similar and yields exactly the same result.
@@ Appendix C Formulas for GW density
[…]
$\Omega_{\rm sw.}h^{2}(f)=2.65\times 10^{-6}(H_{*}\tau_{\rm sw.})\left(\frac{H_{*}}{\beta}\right)v_{w}\left(\frac{\kappa_{\nu}\alpha}{1+\alpha}\right)^{2}\left(\frac{g_{*}}{100}\right)^{-1/3}\left(\frac{f}{f_{\rm sw.}}\right)^{3}\left(\frac{7}{4+3\left(f/f_{\rm sw.}\right)^{2}}\right)^{7/2}\hskip-17.07182pt,$ (114)
where $\tau_{\rm sw.}=\min\!\left[H_{*}^{-1},\,R_{*}/\bar{U}_{f}\right]$ characterizes the effective duration of the sound-wave source Ellis:2019oqb . Here $R_{*}$ denotes the mean bubble separation, satisfying $H_{*}R_{*}=v_{w}(8\pi)^{1/3}(\beta/H_{*})^{-1}$, and the root-mean-square fluid velocity can be approximated as
$\bar{U}_{f}^{2}\simeq\frac{3}{4}\left(\frac{\kappa_{\nu}\alpha}{1+\alpha}\right).$ (115)
The efficiency factor $\kappa_{\nu}$ is approximated by Espinosa:2010hh  [^8] WallGo defines the efficiency factor as $\kappa_{\rm WallGo}=E_{\rm kin}/\Delta\rho_{\rm vac}$, which is related to $(\kappa_{\nu}\alpha)/(1+\alpha)=\kappa_{\rm WallGo}\alpha/(1+\alpha)$. Throughout this work we use the WallGo definition, which agrees well with Eq. (116) over most of the parameter space we explore.
$\kappa_{\nu}\simeq\begin{cases}\alpha\left(0.73+0.83\sqrt{\alpha}+\alpha\right)^{-1},&v_{w}\sim 1,\\ 6.9\,v_{w}^{6/5}\alpha\left(1.36-0.037\sqrt{\alpha}+\alpha\right)^{-1},&v_{w}\ll 1.\end{cases}$ (116)
The corresponding peak frequency is
$f_{\rm sw.}=1.9\times 10^{-5}\,\frac{1}{v_{w}}\left(\frac{\beta}{H_{*}}\right)\left(\frac{T_{*}}{100\,{\rm GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6}{\rm Hz},$ (117)
where $g_{*}$ denotes the number of relativistic degrees of freedom.
On the other hand, magnetohydrodynamic turbulence in the plasma provides a subleading contribution to the gravitational-wave signal, with energy density
$\Omega_{\rm turb.}h^{2}(f)=3.35\times 10^{-4}\left(\frac{\beta}{H_{*}}\right)^{-1}\left(\frac{\epsilon\,\kappa_{\nu}\alpha}{1+\alpha}\right)^{3/2}\left(\frac{g_{*}}{100}\right)^{-1/3}v_{w}\frac{(f/f_{\rm turb.})^{3}(1+f/f_{\rm turb.})^{-11/3}}{1+8\pi f/h_{*,0}},$ (118)
where the redshift factor to today is
$h_{*,0}=16.5\left(\frac{T_{*}}{10^{8}\,{\rm GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6}{\rm Hz}.$ (119)
We assume that the turbulence efficiency factor can be written as $\kappa_{\rm turb}=\epsilon\,\kappa_{\nu}$, where $\epsilon$ parametrizes the fraction of bulk fluid motion converted into turbulence. Numerical simulations of magnetohydrodynamic turbulence Hindmarsh:2015qta indicate that at most $5\%$$10\%$ of the kinetic energy in bulk flows is transformed into vorticity. We therefore adopt the conservative choice $\epsilon=0.05$ Caprini:2015zlo , which we employ throughout this work. The peak frequency is
$f_{\rm turb.}=2.7\times 10^{-5}\,\frac{1}{v_{w}}\left(\frac{\beta}{H_{*}}\right)\left(\frac{T_{*}}{100\,{\rm GeV}}\right)\left(\frac{g_{*}}{100}\right)^{1/6}{\rm Hz}.$ (120)
Ref. Caprini:2024gyk recently proposed a new generation of spectral templates encompassing weak, intermediate, and strong first-order phase transitions, with the strong regime ($\alpha\sim 0.5$) constructed for the first time. While the weak-transition limit is well described by earlier results, intermediate and strong transitions exhibit deviations from the conventional linear scaling with the source duration. For strongly supercooled transitions proceeding in a vacuum-energy–dominated background, we adopt the framework of Ref. Caprini:2024hue ,
$h^{2}\Omega_{\rm GW}(f)=h^{2}\Omega_{p}\left(\frac{f}{f_{p}}\right)^{n_{1}}\left[1+\left(\frac{f}{f_{p}}\right)^{n_{1}-n_{2}}\right]^{-1},$ (121)
with slopes
$n_{1}=2.4\quad(f<f_{\rm bub}),\qquad n_{2}=-2.4\quad(f>f_{\rm bub}),$ (122)
where $f_{\rm bub}$ is the peak frequency associated with bubble collisions. The mapping between thermodynamic and spectral quantities is
$h^{2}\Omega_{p}=h^{2}F_{\rm GW,0}\,A_{\rm str}\,\tilde{K}^{2}\left(\frac{H_{*}}{\beta}\right)^{2},$ (123)
$f_{\rm bub}\simeq 0.11\,h_{*,0}\left(\frac{\beta}{H_{*}}\right),$ (124)
@@
$\tilde{K}=\frac{\alpha}{1+\alpha},\qquad A_{\rm str}\simeq 0.05,$ (125)
and
$h^{2}F_{\rm GW,0}\simeq 1.64\times 10^{-5}\left(\frac{100}{g_{*}}\right)^{1/3}.$ (126)
Finally, the bubble-wall velocity $v_{w}$ is determined either macroscopically from hydrodynamic matching or dynamically once plasma friction is included Steinhardt:1981ct ; Moore:1995ua ; John:2000zq ; Espinosa:2010hh ; Bodeker:2009qy . Assuming local thermal equilibrium and an infinitesimally thin wall, the plasma is described by a perfect-fluid energy–momentum tensor $T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+pg^{\mu\nu}$. Energy–momentum conservation across the wall yields the Rankine–Hugoniot conditions
$w_{+}\gamma_{+}^{2}v_{+}$ $=w_{-}\gamma_{-}^{2}v_{-},$ (127)
$w_{+}\gamma_{+}^{2}v_{+}^{2}+p_{+}$ $=w_{-}\gamma_{-}^{2}v_{-}^{2}+p_{-},$ (128)
with $w=\rho+p$. These equations admit deflagration, detonation, and hybrid solutions, but do not uniquely determine $v_{w}$, allowing instead a continuous family of solutions.
Including deviations from equilibrium leads to a friction force acting on the wall. The scalar background $\phi(z)$ satisfies
$\phi^{\prime\prime}-\frac{\partial V_{\rm eff}}{\partial\phi}+\sum_{i}\frac{\partial m_{i}^{2}}{\partial\phi}\int\frac{d^{3}p}{(2\pi)^{3}2E}\,\delta f_{i}=0,$ (129)
where $\delta f_{i}$ obey Boltzmann equations in the wall frame. In the relaxation-time approximation, the resulting friction force is linear in the wall velocity,
$F_{\rm fric}=\eta\,v_{w},\qquad\eta\simeq\sum_{i}\frac{g_{i}}{T}\left(\frac{\partial m_{i}^{2}}{\partial\phi}\right)^{2}\int dz\,\frac{\phi^{\prime}(z)^{2}}{\Gamma_{i}},$ (130)
with $\Gamma_{i}$ the interaction rates. The terminal velocity follows from the force-balance condition $\Delta V_{\rm eff}(T)=\eta v_{w}$. If friction cannot balance the vacuum pressure, the wall accelerates into a runaway regime.
In practice, the bubble-wall velocity is obtained by either solving the hydrodynamic matching equations with a suitable selection criterion or by numerically solving the coupled scalar-field and Boltzmann equations. In this work, we employ either approach depending on the validity conditions of WallGo Ekstedt:2024fyq ; vandeVis:2025plm , which we use for the numerical determination of $v_{w}$.